OpenStudy (courtneyb):

Simplify: i137 + i1003

OpenStudy (hunus):

Is that i^137 + i^1003?

terenzreignz (terenzreignz):

offline. I would guess so though. What's your answer? :)

OpenStudy (hunus):

0

terenzreignz (terenzreignz):

It's correct, but show your process, in case the user returns...

OpenStudy (hunus):

When multiplying i, there is a pattern. i*i= -1 -1*i=-i -i*i=1 Every fourth number is a 1 So if we section the power it's raised to into groups of four we can get (i*i*i*i)=1 * (i*i*i*i) = 1... for i^137 137/4 = 34 with a remainder of 1 This is (i*i*i*i)*(i*i*i*i)*(i*i*i*i)... 34 times which is 1 times itself 34 times. multiplied by that remainder of 1 1*i = i So now we have i + i^1003 1003/4 = 250 with a remainder of 3 This is 1 times itself 250 times multiplied by the three remaining i's which is equal to 1*i*i*i= -i So we have i + (-i) = 0

terenzreignz (terenzreignz):

:) Thanks a bunch :D

OpenStudy (hunus):

Yup :D